OpenStudy (diamond77108):

PLZ HELP WILL MEDAL 21/2x−3/4(2x+5)=3/8

alones (alones):

Alright so simplify \[9x+\frac{−15 }{ 4 }=\frac{ 3 }{ 8 }\]

alones (alones):

And add \(\frac{ 15 }{ 4 }\)to sides \[9x+\frac{ -15 }{ 4 }+\frac{ 15 }{ 4 }=\frac{ 3 }{ 8 }+\frac{ 15 }{ 4 }\] \[9x=\frac{ 33 }{ 8 }\]

OpenStudy (diamond77108):

wait that first fraction is 2 1/2

alones (alones):

So now divide by \(9\) \[\frac{ \frac{ 33 }{ 8} }{ 9 }\]

OpenStudy (diamond77108):

37 /18

OpenStudy (diamond77108):

37 1/8

alones (alones):

Oh waitt so it's separate like this? \[2\frac{ 1}{ 2 }x-\frac{ 3 }{ 4 }(2x+5)\frac{ 3 }{ 8 }\]

OpenStudy (diamond77108):

yeah

OpenStudy (diamond77108):

and it equals 3/8

OpenStudy (diamond77108):

my answers are x=−4 5/8​ x=−3 3/8 ​ x=4 1/8 ​ ​ x=5 3/8 ​

alones (alones):

Okay so if your problem looks like this \(2\frac{ 1}{ 2 }x-\frac{ 3 }{ 4 }(2x+5)=\frac{ 3 }{ 8 }\) Add \(\frac{ 15 }{ 4 }\)to both sides. \[x+\frac{ -15 }{ 4 }+\frac{ 15 }{ 4}=\frac{ 3 }{ 8 }+\frac{ 15 }{ 4 }\] which would equal \[x=\frac{ 33 }{ 8 }\]

alones (alones):

And so as a mixed number it would be \[x=4\frac{ 1 }{ 8 }\]

alones (alones):

@diamond77108 (;